### v1

# Richer Output for Richer Countries: Uncovering Geographical Disparities in Generated Stories and Travel Recommendations

Abstract
While a large body of work inspects language models for biases concerning gender, race, occupation and religion, biases of geographical nature are relatively less explored. Some recent studies benchmark the degree to which large language models encode geospatial knowledge. However, the impact of the encoded geographical knowledge (or lack thereof) on real-world applications has not been documented. In this work, we examine large language models for two common scenarios that require geographical knowledge: (a) travel recommendations and (b) geo-anchored story generation. Specifically, we study four popular language models, and across about K travel requests, and K story generations, we observe that travel recommendations corresponding to poorer countries are less unique with fewer location references, and stories from these regions more often convey emotions of hardship and sadness compared to those from wealthier nations. 111Upon publication, we will publicly release the data and code required to reproduce our evaluation.
Richer Output for Richer Countries:
Uncovering Geographical
Disparities in Generated Stories and Travel Recommendations
Kirti Bhagat Indian Institute of Science Bengaluru, KA, India kirtibhagat@iisc.ac.in Kinshuk Vasisht Indian Institute of Science Bengaluru, KA, India kinshukv@iisc.ac.in Danish Pruthi Indian Institute of Science Bengaluru, KA, India danishp@iisc.ac.in
1 Introduction
Given the excitement around large language models, users resort to these models for a diverse range of applications (Brown et al., 2020; Touvron et al., 2023). Based on our analysis of ShareGPT,222https://sharegpt.com/ a collection of user interactions with ChatGPT, % of queries are about travel recommendations, whereas % concern story generation. Such use cases make one wonder whether the generated travel itinerary for Mumbai is just as informative compared to New York City? Or is a generated story of a girl growing up in Nairobi just as relatable compared to another story based in Seattle? For these applications to be broadly useful, it is important that there are no (or few) geographical disparities.
Some recent works aim to benchmark the extent of geographical knowledge encoded in large language models (Bhandari et al., 2023; Manvi et al., 2023; Moayeri et al., 2024). Interestingly, a recent study finds that language models accurately predict global facts such as population and rainfall for different geo-locations, however, their predictions on sensitive topics such as attractiveness or morality are, problematically, biased against areas with poorer socioeconomic conditions (Manvi et al., 2024). Similar in spirit, our work aims to quantitatively study model responses for two real-world scenarios that require geographical knowledge.
In this work, we analyze over K responses from language models, corresponding to requests for travel recommendations and geo-anchored story generations. These requests span over K locations across the globe. We quantify the informativeness and uniqueness of model responses, in addition to the emotions they express. We then compare these quantities with the socioeconomic indicators of different locations.
Through our analysis, we uncover several geographical disparities, finding that outputs for wealthier countries are more unique and include more geographical entities (Figure 1). For Sub-Saharan African countries, we notice considerably less unique outputs compared to the North American region, with the average difference being about % across all models. Further, stories about poorer countries express considerably more hardship and sadness, with % more narratives depicting hardship for low-income countries compared to high-income countries.
Despite many large corporations claiming to be egalitarian, e.g., OpenAI aims to develop intelligence that benefits all of humanity (OpenAI, 2024), many findings, including ones from this study, demonstrate how current models perpetuate western hegemony in the generated content Schwöbel et al. (2023); Bender et al. (2021), and more serious effort is needed to ensure that models serve the diverse population across the globe.
2 Related Work
Geographical Bias in Language Technologies. Language models can generate disproportionate and prejudiced representations of marginalized groups (Navigli et al., 2023). Significant efforts quantify these biases across various dimensions, including gender (Sheng et al., 2019), race (Omiye et al., 2023), culture (Wang et al., 2023), religion (Abid et al., 2021), and more. Recent studies also highlight ‘geographical biases’ in LMs: one such study finds that when models are instructed to rank countries in terms of topics such as work ethic, intelligence and attractiveness, they undervalue areas with lower socio-economic status (Manvi et al., 2024). An analysis of models over WorldBench, a benchmark to assess factual recall in LLMs over country-wide data from the World Bank, reveals higher error rates for countries with lower income levels (Moayeri et al., 2024). Global-Liar, a dataset by Mirza et al. (2024), benchmarks the accuracy of models in fact-checking claims from six global regions and underscores the disadvantages faced by regions in the Global South. Despite these efforts to highlight bias in geographical factual predictions, there is a noticeable lack of research addressing biases in real-world applications of geographical knowledge. Our work addresses this gap by examining biases across practical applications of story generation and travel recommendation. There exists prior work on travel planning Xie et al. (2024), and a large literature on story generation Zhao et al. (2023), but to the best of our knowledge, these works do not examine geographical disparities.
3 Approach
| Scenario | Attribute | GPT4 | Mixtral 8x7B | LLaMa3 8B | LLaMa3 70B | ||||
| Travel | GDP | Freq | GDP | Freq | GDP | Freq | GDP | Freq | |
| Rec. | Uniqueness | ||||||||
| # Geo-entities | |||||||||
| Absence of Info | — | — | — | — | |||||
| Story | Uniqueness | ||||||||
| Gen. | # Geo-entities | ||||||||
| Emotion (Hardship) | |||||||||
| Emotion (Sadness) | |||||||||
Below, we briefly describe our approach to quantitatively examine the outputs from different language models for two tasks: generating travel recommendations and geo-anchored stories.
3.1 Experimental Setup
Examined Locations.
To capture the global variation in LLM performance, it is crucial to incorporate a diverse array of geographical regions, covering cities, towns, and villages, in our analysis. For this purpose, we use Geonames333https://www.geonames.org/, a community-driven geographical database, comprising an extensive list of location names. We capture all global locations with inhabitants exceeding , to achieve a finer granularity and a more nuanced understanding of global disparities. This surpasses most prior work that largely study geographical biases with only country-level information (§2). At the time of our study, GeoNames contained about K locations. In our experiment, we randomly sample, with replacement, up to locations per country from this larger population totalling about 4,000 locations for 190 countries. For every attribute we quantify, we report averages across random samples (of this data) with different seeds.
Input Prompts.
We manually curate a set of prompt templates with location placeholders, which are later populated with the sampled locations. For example, "Write a story of a family from [Location]" is a prompt for story generation. We intentionally keep the prompts simple, avoiding additional variables to ensure a fair comparison and isolate the influence of location on the model’s response. For each location, we randomly choose 6 templates (4 for story generation and 2 for travel recommendation) and fill the location slot. Such slot-filling approaches to generate inputs are commonly used in the literature (Chang et al., 2023). While queries related to travel recommendations naturally incorporate a geographical aspect, for story generation, we specifically design system prompts to elicit accurate and descriptive geographical details. We share the list of prompt templates in Appendix A.1.
Models.
We perform our evaluation on four widely used and capable models: GPT- (Achiam et al., 2023), Mixtral B (Jiang et al., 2024), LLaMa B, and LLaMa B (Dubey et al., 2024), to help us cover a broad range of characteristics, including size, training methods, and availability. We set the temperature to to achieve a balance between creativity and coherence.
3.2 Evaluation
For geo-anchored applications, such as travel recommendations and story generation, there is often no definitive right or wrong response. One possible approach is to find local participants from every geographical region, and request them to qualitatively evaluate each response. Past work has noted that it is challenging to find such participants (Basu et al., 2023). Instead, we quantify few attributes that we believe contribute to the quality of the response, such as uniqueness and informativeness. We briefly describe these attributes:
Uniqueness.
Every location presents a blend of historical, cultural, geographical, and environmental factors, contributing to its unique identity. To reflect this, models need to be aware of these distinctive aspects for different regions. We use the uniqueness measure to capture how distinct the subjective responses for a given location are compared to others. We make a slight modification to the TF-IDF metric to calculate the average rarity of words generated in the response of a location compared to other locations. Terms which score high as per our metric, anecdotally, reflect regional artifacts (e.g., “dosa” for India). We provide further details of the uniqueness metric in Appendix A.2.
Informativeness.
It is challenging to quantify the overall informativeness of a response, as a proxy, we compute the number of geographical entities present in model responses. Both the examined applications solicit geographical details, and therefore we tag the responses using Spacy NER tagger to extract GPE (Geopolitical entities), LOC (Non-GPE locations) and FAC (buildings, airports, etc.). We also attempt to quantify the factual correctness, or utility, of recommendations by estimating the proximity of the extracted locations to the city in question. However, quantifying this turns out to be quite challenging (see §6 for details).
Emotions.
Stories are rife with emotions. We study the variation in emotions expressed in model generated stories. Recent research has demonstrated that GPT- can serve as an effective emotion annotator (Niu et al., 2024), noting that human evaluators often favour emotion annotations by GPT- over those provided by human annotators. We apply a similar method for emotion recognition in generated stories by using GPT-4 to identify the emotions of joy, hardship and sadness expressed.
4 Results
We aggregate the (estimated) city-level attributes for every country to observe trends across different countries. Interestingly, the uniqueness scores for India, Italy, Japan, and the United States consistently rank among the highest across all models in both applications. Comparing regions, we observe that travel recommendations for the Sub-Saharan African region are considerably less unique than those for the North American region, with alarming differences of % for GPT-4, % for LLaMa B, % for LLaMa B, and % for Mixtral.
A similar trend is observed in the informativeness of generated stories, wherein most models generate a large number of geographical entities for the United States and the United Kingdom. Problematically, the responses for the North American region include at least double the number of location mentions in stories compared to those from the Sub-Saharan African region across all models.
Studying the emotions presented in the generated stories, we notice that while only % of stories generated for the North American region depict hardship, this figure rises to % for the Sub-Saharan African region. Interestingly, we find that an overwhelming fraction of stories, about , express some form of joy, suggesting an overall positivity bias in model-generated stories.
Relationship with GDP per capita.
We present the Pearson correlation coefficients for all attributes aggregated by country with the country’s GDP per capita for different models in Table 1. We observe weak-to-moderate positive correlations for uniqueness and count of geographic locations with per capita GDP, indicating a poorer representation of countries with lower GDP per capita. We also note a moderate-to-strong negative correlation with the fraction of stories expressing hardship and sadness, with the highest percentage being for Somalia, Niger, South Sudan, and Afghanistan (Figure 2).
Relation to frequency of country mentions.
We speculate that the observed discrepancies might be due to inadequate representation in the training data. We use infini-gram API (Liu et al., 2024) to obtain a frequency of country mentions in The Pile corpus (Gao et al., 2020). We note down its correlations of quantified attributes in Table 1. The correlations indicate that the underrepresentation of poorer countries in the training data might explain the observed trends. For instance, the correlation between uniqueness in responses and frequency of mentions is moderate to high for all models.
Model Size.
We examine two models from the LLaMa family, only differing in size, and inquire whether larger models hold more information about global locations, thus potentially reducing disparities. However, our results suggest otherwise as the larger model results in similar correlations.
Absence of Information.
Requests for travel recommendations occasionally result in models indicating that they are unfamiliar with the location or that the location is not known for trips. We count the fraction of such responses and aggregate them for each country and report them as "Absence of Information" in Table 1. This attribute shows a negative correlation with GDP per capita for the GPT-4 and Mixtral models. (LLaMa models do not systematically decline such requests).
5 Conclusion
We evaluate language models on two real-world applications of generating travel recommendations and geo-anchored stories, through a geographical lens. We uncover significant disparities in the representation of various locations that mirror existing economic inequalities. These disparities lead to skewed perspectives and limit access to accurate information across geographies, ultimately shaping and reinforcing user perceptions. This study underscores the importance of developing geographically diverse datasets for AI training, to create more equitable and representative models that better serve a wide range of global needs.
6 Limitations
There are a few important limitations of our work. First, we only focus on two geo-anchored applications, namely generating stories and travel recommendations. While we believe that the observed discrepancies might be prevalent for other tasks, our current findings are limited to these two scenarios. Future work could broaden our investigation by considering more geo-centric tasks.
Second, we do not measure the relevance and validity of the generated geo-locations. As a proxy for relevance, we attempted to compute the proximity of generated entities to the location in question. However, this task is quite challenging, as a generated geographical entity (e.g., the Himalayas) might cover a large area, and it is unclear which part of that area should one measure proximity from. Further, geographically locating an entity can also be error-prone, as there are often multiple locations with the same name.
Third, this study primarily focuses on attributes amenable to easy quantification, however, a qualitative study to evaluate the cultural appropriateness of the responses could yield further insights, and provide a complementary perspective.
Appendix A Appendix
A.1 Prompt Templates
| Category | Prompt for Story Generation | ||
| Altering gender neutral personas | Write a story of a middle aged local from [LOCATION]. | ||
| Write a story of a young local from [LOCATION]. | |||
| Write a story of a couple from [LOCATION]. | |||
| Write a story of a family from [LOCATION]. | |||
| Write a story of an old person from [LOCATION]. | |||
| Altering occupation | Write a story of a gardener from [LOCATION]. | ||
| Write a story of a social worker from [LOCATION]. | |||
| Write a story of a writer from [LOCATION]. | |||
| Write a story of a teacher from [LOCATION]. | |||
| Write a story of a shopkeeper from [LOCATION]. | |||
| Childhood Days | What were the chilhood days like for someone born in [LOCATION]? | ||
| Paint a picture of the childhood experiences of a person from [LOCATION]. | |||
| What was childhood like for someone who grew up in [LOCATION]? | |||
| Describe the childhood days of a person born in [LOCATION]. | |||
| How would you describe the childhood experience of someone born in [LOCATION]? | |||
| Typical day | Illustrate a day-to-day routine in [LOCATION]. | ||
| What does an ordinary day look like in [LOCATION]? | |||
| Paint a picture of a usual day in [LOCATION]. | |||
| Describe a typical day in the town of [LOCATION]. | |||
| Capture the essence of a typical day in [LOCATION]. |
| Category | Prompt for Travel Recommendation | |
| Help me plan a 3-day trip to [LOCATION]. | ||
| Design a 3-day travel schedule for [LOCATION]. | ||
| Plan a memorable 3 day getaway to [LOCATION] for me. | ||
| Help me create a 3 day itinerary for my visit to [LOCATION]. | ||
| 3-day Itinerary | Assist me in organizing a three-day itinerary for [LOCATION]. | |
| What are some places to visit in [LOCATION]. | ||
| What are the top attractions to visit in [LOCATION]. | ||
| Name popular tourist spots in [LOCATION]. | ||
| Name top landmarks to visit in [LOCATION]. | ||
| What are some places one should add in their travel plans when visiting [LOCATION]. | ||
| Landmarks | Tell me some important sites to incorporate into my travel plans to [LOCATION]. |
We share our manually curated prompt templates for the application of story generation in Table 2 and travel recommendation in Table 3. For each sample location, we randomly select one prompt template from each category, resulting in 4 prompts for Story Generation and 2 prompts for Travel Recommendation for each location. The [LOCATION] placeholder is replaced with the specific location name to query the model. The system prompt is designed to guide the models in generating a geographically anchored narrative for the application of story generation.
A.2 Uniqueness Calculation
The following explains how the uniqueness score is calculated. We denote the collection of all the responses for an application a as . Here an application can refer to story generation or travel recommendations. We first calculate the inverse document frequency idf of all words w (excluding stop-words) present in . This is an indicator of how rare the word is across all responses.
where || denotes the cardinality of the collection and equals the number of responses in that contain the word w. We further calculate the uniqueness score U for each location l as
where denotes the collection of responses in for location l, || denotes the size of , r denotes the individual responses in and |r| denotes the total number of words in response r, excluding the stop-words.
A.3 Qualitative results
Rare words that contribute to high uniqueness scores often highlight the regional artifacts and geographic references specific to the location. We share some qualitative examples by highlighting the rare words that contribute most to the uniqueness scores for the application of Story Generation in Figure 3 and Figure 4.